OpenStudy (anonymous):

find the value of x such as y' of y=x^2/3 - x^1/3 is 1/4 a. 9 b. 10 c. 8 d. 11

OpenStudy (anonymous):

do you think the value of x is 8?

OpenStudy (anonymous):

i can get it w8

OpenStudy (radar):

(2/3)(1/x^(1/3)-(1/3)(1/x^2/3) substitution 8 for x 2/6-1/12=3/12=1/4 yes, x=8

OpenStudy (anonymous):

theres something wrong w8

OpenStudy (anonymous):

if 8 for x i trololo it but theres something tricky i cant see

OpenStudy (radar):

Actually I never was able to solve x for 8, but just substituted 8 for x and it works out.

OpenStudy (anonymous):

oww ok ok